OpenStudy (anonymous):

Find the equation of the circle whose center and radius are given. center ( 5, 6), radius = 3

OpenStudy (radar):

(x-5)^2 + (y-6)^2 = (3)^2 or 9

OpenStudy (anonymous):

(x-h)^2+(y-k)^2=r^2,h=5.k=6 and r=3 x^2+y^2-10x-12y+42=0,is the equation

OpenStudy (anonymous):

not 42,its 52

OpenStudy (anonymous):

The question that is to be filled in, = one number

OpenStudy (anonymous):

The equation I need to fill in is (x___________=_

OpenStudy (radar):

Do the math: (x-5)^2=x^2-10x+25 (y-6)^2=y^2-12y+36 I would combine them and get \[x ^{2}-10x+y ^{2}-12y+61=9\] actually I would leave it as: \[(x ^{2}-10x+25) + (y ^{2}-12y + 36) =9\]

OpenStudy (anonymous):

So is it the first equation that you posted that would fill it in or that one?

OpenStudy (radar):

I liked the first one, but they all are correct, I don't know what your instructor is expecting as which form is the "correct" one. I would present the first one, and then work it out to be looking like the last one. I don't know if you have that luxury or not, if it is an on-line course which one do you have as an option?